Determine the values of the node voltages V1 and v2 in this circuit: 30 V 23 Ω Ι 41 Ω V2 41 Ω Ι 41 Ω 15 V 15 vΟ 37 Ως 10 V
Answer #1

assol = ntos = EA ne8t.o-ra nsh.al-an of ③ ubo buinus ho ③ - Istornstoot in 8 ho.or (+)+(- -) t = 11-04-2nt ont Toon into E=